# bHUN-REN Wigner Research Centre for Physics, Konkoly-Thege Miklos ut 29-33, 1121 Budapest, Hungary

### v1

@@ 4.3 Median resummation
This completes the full resurgence structure of the strong coupling expansion of (1) and provides strong evidence on the validity of the transseries structure (52). It would be interesting to investigate how the resurgence structure discussed above manifests itself in the strong coupling expansion of the cusp anomalous dimension given by (8).
@@ Appendix A Bridge equations
and $\lambda(\delta)$ is a function of $\delta$ that governs the leading $1/g$ order at exponential level $\delta$. For simplicity in this Appendix I considered only one sector of exponential weights. Although the large-$g$ expansion of the determinant $Z(g)$ in (52) contains two sets of weights $x^{+}_{j}$ and $x^{-}_{j}$, the argument below remains the same for $Z(g)$ as well. The function $F(g)$ with $\lambda(\delta)=0$ and $g\to 8\pi g$ coincides with the strong coupling expansions investigated in Bajnok et al. (2025d).
@@
$\mathcal{S}_{\pm}\left[\Phi^{(\delta)}\right](g)=g\int_{0}^{\infty e^{i\epsilon}}dse^{-gs}\mathcal{B}\left[\Phi^{(\delta)}\right](s)\,.$ (135)

### v2

@@ 4.3 Median resummation
This completes the full resurgence structure of the strong coupling expansion of (1) and provides strong evidence on the validity of the transseries structure (52). As a physical example, in Bajnok et al. (2026a) we summarize the complete strong coupling and resurgence structure of the cusp anomalous dimension.
@@ Appendix A Bridge equations
and $\lambda(\delta)$ is a function of $\delta$ that governs the leading $1/g$ order at exponential level $\delta$. For simplicity in this Appendix I considered only one sector of exponential weights. Although the large-$g$ expansion of the determinant $Z(g)$ in (52) contains two sets of weights $x^{+}_{j}$ and $x^{-}_{j}$, the argument below remains the same for $Z(g)$ as well. The function $F(g)$ with $\lambda(\delta)=0$ and $g\to 8\pi g$ coincides with the strong coupling expansions investigated in Bajnok et al. (2026b).
@@
$\mathcal{S}_{\pm}\left[\Phi^{(\delta)}\right](g)=g\int_{0}^{\infty e^{\pm i\epsilon}}dse^{-gs}\mathcal{B}\left[\Phi^{(\delta)}\right](s)\,.$ (134)